a gas occupies 40 ml at -123 degrees celcius.What volume does it occupy at 27 degrees celcuis?
1
Expert's answer
2020-05-06T13:57:54-0400

"applying \\\frac{V_1}{T_1}=\frac{V_2}{T_2}"

"\frac{40}{273-123}=\frac{V_2}{273+27}"


"V_2=40\times300/150=80 ml"